In a class of 30 students the ratio of boys to girls is 3:7. If 8 girls join the class find the ratio of girls to boys in the class
Answer by TimothyLamb(4379)

x = number of boys
y = number of girls
---
x + y = 30
x/y = 3/7
7x = 3y
---
put the system of linear equations into standard form
---
x + y = 30
7x - 3y = 0
---
copy and paste the above standard form linear equations in to this solver:
https://sooeet.com/math/system-of-linear-equations-solver.php
---
solution:
x = number of boys = 9
y = number of girls = 21
---
now, add 8 girls:
---
x = number of boys = 9
y = number of girls = 29
---
answer:
y/x = ratio of girls to boys = 29/9
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations with quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php
---
Send and receive free text messages online:
https://sooeet.com